A charge of 2.88 $\mu C$ is held fixed at the origin. A second charge of 2.88 $\mu C$ is released from rest at the position (1.01 m, 0.546 m). If the mass of the second charge is 2.52 g,

(1) what is its speed when it moves infinitely far from the origin?

(2) At what distance from the origin does the 2.88 $\mu C$ charge attain half the speed it will have at infinity?
Answer #1

q1 = 2.88*10^-6 C


(x1,y1) = (0,0)


q2 = 2.88*10^-19 C


(x2,y2) = (1.01 , 0.546)

distance between q1 and q2 , r1 = sqrt((x2-x1)^2 + (y2-y1)^2)

r1 = sqrt(1.01^2+0.546^2)


r1 = 1.15 m

initial potential energy Ui = k*q1*q2/r1

initial kinetic energy (1/2)*m*vi^2 = 0


at infinity

rf = infinity

final potential energy Uf = k*q1*Q2/rf = 0

final kinetic energy Kf = (1/2)*m*vf^2

from energy conservation


total energy is constant

Ef = Ei

Uf + Kf = Ui + Ki

0 + (1/2)*m*vf^2 = k*q1*q2/r + 0


(1/2)*2.52*10^-3*vf^2 = 9*10^9*2.88*10^-6*2.88*10^-6/1.15


speed vf = 7.18 m/s

===================================================


part b

at distance r2 the speed v2 = vf/2 = 7.18/2 = 3.59 m/s


initial potential energy Ui = k*q1*q2/r1

initial kinetic energy (1/2)*m*vi^2 = 0

at distance r2


final potential energy Uf = k*q1*Q2/r2

final kinetic energy Kf = (1/2)*m*v2^2

from energy conservation


total energy is constant

Ef = Ei

Uf + Kf = Ui + Ki

k*q1*q2/r2 + (1/2)*m*v2^2 = k*q1*q2/r1 + 0


9*10^9*2.88*10^-6*2.88*10^-6/r2 + (1/2)*2.52*10^-3*3.59^2 = 9*10^9*2.88*10^-6*2.88*10^-6/1.15

r2 = 1.534 m <<<-----------ANSWER